Every real valued function can be expressed uniquely as the sum of an even and odd function.


Is this true..? if yes please explain with some example.!

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Mukul

take a function y=f(x)

f(x)can be written as

f(x)= [f(x)+f(-x)]/2  +   [f(x)-f(-x)]/2

     = h(x)               +   g(x)

h(x)= [f(x)+f(-x)]/2

h(-x)=[f(-x)+f(x)]/2 = h(x)  ----->Even function

now g(x)=[f(x)-f(-x)]/2

  g(-x)=[f(-x)-f(x)]/2  = -[f(x)-f(-x)]/2=-g(x)-------> ODD function

hence every real valued function f(x) can be expressed as a sum of function h(x) and g(x) which are even and odd respectively. you can verify it by taking f(x) some function. try it.

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..